Solve the equation.
z/4 = 11/10
Select the correct answer.
a. z = 1/5
b. z = 2/3
c. z = 24
d. z = 22/5
*note the equation and answers are fractions

Answer by richwmiller(17219)

z/4=11/10
z=44/10
d. z = 22/5